### v3

# Pabitra Barik

Abstract
Let $C$ be a smooth complex projective curve of genus $g>1$ and $A=J(C)$ its Jacobian with principal polarization $\Theta$. Starting from a semistable vector bundle $V$ on $C$ with $\mu(V)>2g-2$, we consider the Fourier–Mukai transform $E=\Phi_{\mathcal{P}}(a_{*}V)$. We prove that $E(\Theta)$ satisfies the $\mathrm{IT}_{0}$ property.
1 Introduction
Let $C$ be a smooth projective curve of genus $g\geq 2$, and let $A=J(C)$ denote its Jacobian equipped with the principal polarization $\Theta$. Our approach is based on Fourier–Mukai techniques and exploits the geometry of the Abel–Jacobi embedding $C\hookrightarrow J(C)$. Starting from a sufficiently positive semistable vector bundle bundle $V$ on $C$, we consider the pushforward $a_{*}V$ along the Abel–Jacobi map and study its Fourier–Mukai transform. After a suitable twist by the principal polarization, we show that the resulting bundle on $J(C)$ satisfies strong regularity properties in the sense of Pareschi–Popa [4].
We recall some preliminaries.
1.1 The Abel–Jacobi embedding
Fix a base point $x_{0}\in C$. The Abel–Jacobi morphism
$a:C\longrightarrow A,\qquad x\longmapsto\mathcal{O}_{C}(x-x_{0}),$
is a closed immersion. We identify $C$ with its image in AA.
A classical result (see [2, Ch. VII]) states that
$a^{*}\Theta\sim\text{a divisor of degree }g\text{ on }C.$
In particular,
$\deg(a^{*}\mathcal{O}_{A}(\Theta))=g.$
1.2 Fourier–Mukai transform on the Jacobian
Let $\widehat{A}:=Pic^{0}(A)$ be the dual abelian variety. Since AA is principally polarized, we identify
$A\cong\widehat{A}$
via $\Theta$.
Let
$\mathcal{P}$
be the normalized Poincaré bundle on $A\times\widehat{A}$, characterized by the properties:
  1. 1.
    $\mathcal{P}|_{A\times\{0\}}\cong\mathcal{O}_{A}$,
  2. 2.
    $\mathcal{P}|_{\{0\}\times\widehat{A}}\cong\mathcal{O}_{\widehat{A}}$.
Let $p_{1}$ and $p_{2}$ denote the projections from $A\times\widehat{A}$ to the first and second factor. The Fourier–Mukai functor
$Db(A^)\Phi_{\mathcal{P}}$$\longrightarrow D^{b}(\widehat{A})$
is defined by
$\Phi_{\mathcal{P}}(\mathcal{F}):=Rp_{2*}(p_{1}^{*}\mathcal{F}\otimes\mathcal{P}).$
1.3 WIT and IT conditions
Let $\mathcal{F}$ be a coherent sheaf on AA.
Definition $1.$
We say that $\mathcal{F}$ satisfies WITi (Weak Index Theorem of index ii) if
$R^{j}\Phi_{\mathcal{P}}(\mathcal{F})=0\quad\text{for all }j\neq i.$
If $\mathcal{F}$ satisfies WITi, we denote
$\widehat{\mathcal{F}}:=R^{i}\Phi_{\mathcal{P}}(\mathcal{F}).$
Definition 1.2.
A coherent sheaf $\mathcal{E}$ on AA satisfies IT0 (Index Theorem with index 0) if
$H^{i}(A,\mathcal{E}\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in Pic^{0}(A).$
IT0 is stronger than WIT0. If $\mathcal{E}$ is locally free and satisfies IT0, then $\mathcal{E}$ is continuously globally generated in the sense of [4].
1.4 Base change for Fourier–Mukai
Let $\mathcal{F}$ be a coherent sheaf on AA. For $\alpha\in Pic^{0}(A)$, derived base change gives
$\Phi_{\mathcal{P}}(\mathcal{F})\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(A,\mathcal{F}\otimes\mathcal{P}_{\alpha}).$
This identity will be used repeatedly to compute fibers of Fourier–Mukai transforms.
The preceding notions provide the foundational framework for the construction carried out in the subsequent sections.
1.5 Slope Stability
Let $(X,H)$ be a smooth projective variety of dimension $n$ with polarization $H$. For a torsion-free coherent sheaf $\mathcal{E}$ on $X$, the slope with respect to $H$ is
$\mu_{H}(\mathcal{E}):=\frac{c_{1}(\mathcal{E})\cdot H^{n-1}}{\operatorname{rk}(\mathcal{E})}.$
Definition 1.3.
  1. 1.
    The sheaf $\mathcal{E}$ is called $\mu_{H}$-stable if for every proper nonzero subsheaf $\mathcal{F}\subset\mathcal{E}$ with $0<\operatorname{rk}(\mathcal{F})<\operatorname{rk}(\mathcal{E})$, one has
    $\mu_{H}(\mathcal{F})<\mu_{H}(\mathcal{E}).$
  2. 2.
    The sheaf $\mathcal{E}$ is called $\mu_{H}$-semistable if the above inequality is replaced by
    $\mu_{H}(\mathcal{F})\leq\mu_{H}(\mathcal{E}).$
  3. 3.
    The sheaf $\mathcal{E}$ is called $\mu_{H}$-polystable if it is a direct sum of $\mu_{H}$-stable sheaves all having the same slope.
(See [7, Ch. 1] )
In particular, a direct sum of $\mu_{H}$-stable sheaves of identical slope is $\mu_{H}$-polystable. For example, on an abelian variety, a direct sum of algebraically trivial line bundles is $\mu_{H}$-polystable of slope zero.
Curves. If $C$ is a smooth projective curve, then
$\mu(V)=\frac{\deg(V)}{\operatorname{rk}(V)}$
for any vector bundle $V$ on $C$, and semistability is independent of a choice of polarization. Semistability is preserved under tensoring with line bundles of degree zero.
2 Cohomological Vanishing and the WIT0 Property
In this section we analyze the Fourier–Mukai transform of the pushforward $a_{*}V$ of a semistable vector bundle $V$ on $C$. The goal is to establish uniform cohomological vanishing on $C$, derive the WIT0 property for $a_{*}V$, and prove that its Fourier–Mukai transform is locally free with explicitly computable rank.
2.1 Slope positivity and uniform $H^{1}$-vanishing
Let $V$ be a vector bundle on $C$. Recall that $V$ is semistable if
$\mu(W)\leq\mu(V)\quad\text{for every proper subbundle }W\subset V.$
We assume throughout this section that
$\mu(V)>2g-2.$
Lemma 2.1 (Uniform $H^{1}$-vanishing).
If $V$ is semistable and $\mu(V)>2g-2$, then
$H^{1}(C,V\otimes M)=0\quad\text{for all }M\in Pic^{0}(C).$
Proof.
For $M\in Pic^{0}(C)$, the bundle $V\otimes M$ remains semistable and satisfies
$\mu(V\otimes M)=\mu(V).$
By Serre duality,
$H^{1}(C,V\otimes M)\cong\operatorname{Hom}(V\otimes M,\omega_{C})^{\vee}.$
Since $\omega_{C}$ is a line bundle of slope 2g22g-2, and $\mu(V)>2g-2$, semistability implies
$\operatorname{Hom}(V\otimes M,\omega_{C})=0.$
The claim follows. ∎
Remark $2.$
The slope inequality $\mu(V)>2g-2$ is sharp for this argument. It ensures that $V$ admits no nonzero morphism to $\omega_{C}$, which is precisely the obstruction to the vanishing of $H^{1}$.
2.2 Derived fibers and WIT0
Let
$E:=\Phi_{\mathcal{P}}(a_{*}V).$
We now compute the derived fiber of $E$ at a point $\alpha\in Pic^{0}(A)$.
By derived base change for the Fourier–Mukai transform,
$E\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(A,a_{*}V\otimes\mathcal{P}_{\alpha}).$
Using the projection formula and $R\Gamma(A,a_{*}(-))\cong R\Gamma(C,-)$, we obtain
$E\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(C,V\otimes a^{*}\mathcal{P}_{\alpha}).$
Since $a^{*}\mathcal{P}_{\alpha}\in Pic^{0}(C)$, Lemma 2.1 implies that
$H^{1}(C,V\otimes a^{*}\mathcal{P}_{\alpha})=0\quad\forall\alpha.$
Thus the derived fiber is concentrated in degree 0.
Proposition 2.3 (WIT0 and local freeness).
Assume $\mu(V)>2g-2$. Then:
  1. 1.
    $a_{*}V$ satisfies WIT0 with respect to $\Phi_{\mathcal{P}}$;
  2. 2.
    $E$ is a locally free sheaf on AA;
  3. 3.
    $\operatorname{rk}(E)=h^{0}(C,V)$.
Proof.
Since for every $\alpha$ the derived fiber $E\otimes^{\mathbf{L}}k(\alpha)$ is concentrated in degree 0, we conclude that
$R^{i}\Phi_{\mathcal{P}}(a_{*}V)=0\quad\text{for all }i>0.$
Thus $a_{*}V$ satisfies WIT0.
Moreover,
$E\otimes k(\alpha)\simeq H^{0}(C,V\otimes a^{*}\mathcal{P}_{\alpha}),$
and the dimension of this space is constant in $\alpha$ by semicontinuity. Hence $E$ is locally free.
Finally,
$\operatorname{rk}(E)=h^{0}(C,V).$
2.3 Rank formula
Applying Riemann–Roch on $C$ and using Lemma 2.1, we obtain:
Corollary 2.4 (Rank formula).
If $\mu(V)>2g-2$, then
$\operatorname{rk}(E)=\deg(V)+\operatorname{rk}(V)(1-g).$
Proof.
Since $H^{1}(C,V)=0$, Riemann–Roch gives
$h^{0}(C,V)=\deg(V)+\operatorname{rk}(V)(1-g).$
The result follows from Proposition 2.3. ∎
The results of this section show that positivity of slope on the curve translates into strong structural properties of the Fourier–Mukai transform on the Jacobian. In particular, $E$ is a well-behaved vector bundle whose rank and fibers admit explicit description.
The next section upgrades this WIT0 property to an IT0 statement by a direct cohomological reduction to the curve.
3 From WIT0 to IT0
In Section 3 we proved that
$E:=\Phi_{\mathcal{P}}(a_{*}V)$
is locally free and that $a_{*}V$ satisfies WIT0. In order to construct Ulrich bundles, we now establish directly that a suitable twist of $E$ satisfies IT0.
We prove the required cohomological vanishing by an explicit reduction to the curve.
4.1 Reduction to cohomology on the curve
Let $\alpha\in\mathrm{Pic}^{0}(A)$. By definition of the Fourier–Mukai transform and derived base change (see Mukai [1, 3]), we have
$E\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(A,a_{*}V\otimes\mathcal{P}_{\alpha}).$
Using the projection formula and the identity $R\Gamma(A,a_{*}(-))\simeq R\Gamma(C,-)$, we obtain
$E\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(C,V\otimes a^{*}\mathcal{P}_{\alpha}).$
Twisting by $\Theta$ and taking cohomology yields
$H^{i}(A,E(\Theta)\otimes\alpha)\simeq H^{i}(C,V\otimes a^{*}\mathcal{O}_{A}(\Theta)\otimes a^{*}\mathcal{P}_{\alpha}).$
Since $a^{*}\mathcal{P}_{\alpha}\in\mathrm{Pic}^{0}(C)$, this reduces the IT0 condition for $E(\Theta)$ to uniform vanishing statements on the curve.
4.2 Slope computation
Recall that
$\deg(a^{*}\mathcal{O}_{A}(\Theta))=g.$
Hence
$\mu\!\left(V\otimes a^{*}\mathcal{O}_{A}(\Theta)\right)=\mu(V)+g.$
Under the standing assumption
$\mu(V)>2g-2,$
we obtain
$\mu(V)+g>3g-2>2g-2.$
Therefore the bundle
$V\otimes a^{*}\mathcal{O}_{A}(\Theta)$
is semistable of slope strictly greater than 2g22g-2.
4.3 Vanishing of $H^{1}$
By Lemma 3.1 applied to
$V\otimes a^{*}\mathcal{O}_{A}(\Theta),$
we obtain
$H^{1}\!\left(C,V\otimes a^{*}\mathcal{O}_{A}(\Theta)\otimes M\right)=0\quad\forall\,M\in\mathrm{Pic}^{0}(C).$
In particular,
$H^{1}(A,E(\Theta)\otimes\alpha)=0\quad\forall\alpha\in\mathrm{Pic}^{0}(A).$
4.4 Vanishing of higher cohomology
Since $C$ has dimension 11, we have
$H^{i}(C,-)=0\quad\text{for }i\geq 2.$
Therefore
$H^{i}(A,E(\Theta)\otimes\alpha)=0\quad\forall i\geq 2.$
4.5 Conclusion
Combining the above vanishing results, we obtain:
Theorem 3.1 (IT0 property of $E(\Theta)$).
Let $V$ be a semistable vector bundle on $C$ with $\mu(V)>2g-2$. Then
$E(\Theta)$
satisfies IT0, i.e.
$H^{i}(A,E(\Theta)\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in\mathrm{Pic}^{0}(A).$
Proof.
The vanishing for $i=1$ follows from the slope computation and Lemma 3.1. The vanishing for $i\geq 2$ follows from the fact that cohomology on the curve vanishes in degrees $\geq 2$. ∎